Length of rectangle is 120 inches more than the width and the perimeter is 892 inches. What are the dimensions of the rectangle?
Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length = 120 + x
Perimeter of rectangle
= 2(Width) + 2(Length)
892 = 2(x) + 2(120 + x)
892 = 2x + 240 + 2x
Collect like terms:
2x + 2x = 892 - 240
4x = 652
x = 163
Width = 163 inches
Length = 283 inches
Hope this helps. :-)